# No time to train! Training-Free Reference-Based Instance Segmentation

### v2

@@ 5 Conclusions
In this work, we introduced a novel training-free approach for few-shot instance segmentation by integrating SAM’s mask generation capabilities with the fine-grained semantic understanding of DINOv2. Our method uses reference images to construct a memory bank, refines its internal representations with feature aggregation, and performs feature matching for novel instances using cosine similarity and semantic-aware soft merging. We demonstrate that careful engineering of existing frozen foundation models can lead to state-of-the-art performance without the need for additional training: we achieve 36.8% nAP on COCO-FSOD (outperforming fine-tuned methods), 71.2% nAP50 on PASCAL VOC Few-Shot, and strong generalisation across domains (as evidenced on CD-FSOD benchmark). Furthermore, our semantic segmentation results confirm that our approach can be extended beyond instance segmentation by aggregating instance predictions into semantic maps.
@@
6 Acknowledgements
Funding for this research is provided in part by the SENSE - Centre for Satellite Data in Environmental Science CDT studentship (NE/T00939X/1), and an EPSRC New Investigator Award (EP/X020703/1). This work used JASMIN, the UK’s collaborative data analysis environment https://jasmin.ac.uk [36].
Supplementary Material
7 Ablation on aggregation strategies
@@
The improvements of our soft-merging semantic-aware strategy are shown in Tab. 5. Other aggregation variants, such as covariance similarity, instance softmax, score decay, iterative mask refine, attention-guided global average, underperformed.
@@
8 Ablation on efficiency
Our training-free stages are optimised and lightweight. (1) Memory bank construction (0.1 s/img) is computed once, only requiring to encode $n$ reference images per class with DINOv2. Reference-image features are pre-cached for all following steps. (2) Semantic matching (0.0003 s/img) is a fully parallelised dot product of cosine similarities. (3) Soft-merging (0.006 s/img) uses a parallel implementation of NMS. Tab. 6 shows our method yields significant performance gains compared to Matcher [46] and speeds up SAM default automatic mask generator (AMG) by ×$\times 3$ via efficient point sampling, faster mask filtering, and removal of unnecessary post-processing.
@@ 8 Ablation on efficiency
9 Cross-Domain Few-Shot Object Detection
We provide additional visualisations of the six target datasets in the CD-FSOD benchmark. These datasets span diverse and challenging domains, including photorealistic, cartoon, aerial, underwater, and industrial imagery, each presenting unique distribution shifts. Despite these variations, our method achieves strong performance across all domains without any fine-tuning, demonstrating its remarkable cross-domain generalization. Figure 5 shows an overview of the results for the 6 datasets. Figures 6, 7, 9, 8, 10, 11 display more detailed results for each of the datsets. All results are shown for 5-shot setting, using 5 reference images per category.

### v3

@@ 5 Conclusion
In this work, we introduce a novel training-free approach for few-shot instance segmentation by integrating SAM’s mask generation capabilities with the fine-grained semantic understanding of DINOv2. Our method uses reference images to construct a memory bank, refines its internal representations with feature aggregation, and performs feature matching for novel instances using cosine similarity and semantic-aware soft merging. We demonstrate that careful engineering of existing frozen foundation models can lead to state-of-the-art performance without the need for additional training: we achieve 36.8% nAP on COCO-FSOD (outperforming fine-tuned methods), 71.2% nAP50 on PASCAL VOC Few-Shot, and strong generalisation across domains (CD-FSOD benchmark). Furthermore, our semantic segmentation results show that our approach can be extended beyond instance segmentation by aggregating instance predictions into semantic maps.
@@
LLM Usage
Within the scope of this paper, LLMs are used only to aid and polish writing.
Impact Statement
This work aims to advance the field of machine learning. While it may have broader societal implications, we do not identify any specific impacts that require explicit discussion.
Appendix A Appendix
A.1 Reference-Image Ablation
To understand why different reference images lead to performance variation, we observe that randomly sampled reference sets (via different seeds) yield different performances, particularly in low-shot settings. We analyse the statistical properties of COCO novel classes annotations, evaluate per-class reference-image sensitivity, and design simple heuristics for selecting higher-scoring references.
A.1.1 Variance in reference set.
Using different randomly sampled reference images leads to variations in results, as performance depends on the quality of the selected reference images for each class. To quantify this variance, we evaluate our method on the COCO-20i few-shot object detection benchmark using different random seeds to select reference images. Figure 7 shows the standard deviation (std) across 10 runs. We observe that increasing the number of reference images (higher n-shots) reduces result variance, with lower std values. In 1, 2, and 3-shot settings, reference image selection has a more noticeable impact, while for 5 or more shots, the low std demonstrates robustness to reference set variation. This suggests that particularly in low-shot settings, performance is sensitive to the choice of individual reference images.
Figure 7: Variance (mAP) n-shot on COCO-20i. Error bars show std over 10 runs with different reference image sets. Higher variance at low n-shot reflects sensitivity to reference images; it decreases as n increases, demonstrating our method’s robustness.
A.1.2 Reference Masks Analysis
We study three annotation characteristics that intuitively affect prototype quality: (1) mask area (object size), (2) mask center location, and (3) distance to image edges.
Mask area. Figure 8 shows the distribution of mask areas. The distribution is skewed toward small masks, confirming that many annotations contain limited visual detail.
Figure 8: Mask-area distribution for the 20 COCO novel classes. Most annotations are small or medium-sized, as shown by the positively skewed distribution (log-scale y-axis).
Mask center location. Figure 10 shows 2D heatmaps of bounding-box centers. Most classes exhibit a centered bias, while some (e.g., car, chair) are more distributed along the horizontal axis.
Mask distance to edge. Figure 9 shows distance-to-edge histograms. A fraction of annotations overlap with image boundaries, implying that the object is partially cropped.
Figure 9: Distance from each mask to the four image boundaries. A substantial number of annotations lie close to image edges, indicating possible partial cropping or truncated objects.
Figure 10: 2D heatmaps of mask centers across COCO novel classes. Most annotations are near the image center, though some classes show strong horizontal spread (e.g., car, chair).
Figure 11: Reference-image performance vs. mask area for all COCO novel classes. Each point represents one reference image evaluated on a class-specific reduced validation subset. Larger mask areas consistently correlate with higher scores. Right: visual examples comparing a high-performing and a low-performing reference for two classes. Best viewed when zoomed in.
A.1.3 Heuristics for Reference Selection
In order to identify common patterns linked to higher performance, we sample 100 diverse reference images per class, explicitly covering a range of mask sizes, centers, and edge distances. Each reference is evaluated on a fixed reduced validation subset.
Figure 11 shows per-class scatter plots of mask area vs. performance. All classes show a clear positive trend: larger masks consistently produce stronger reference prototypes which result in higher downstream scores.
Based on these observations we define simple selection heuristics:
  • Area category (based on class-specific quartiles): large ($\geq$75th), medium (25th–50th), small ($\leq$25th).
  • Centeredness: mask center within $\pm$10% of the image center
  • Edge avoidance: mask must be at least $d$ pixels away from any image boundary.
Figure 12: Effect of mask area (left) and centeredness (right) on downstream performance. Medium/large masks significantly outperform small masks, and centered references outperform off-center ones.
Figure 13: 2D score maps of performance as a function of mask-center location. Both the discrete and smoothed KDE versions show higher performance when reference masks lie near the image center.
Figure 12 and Figure 13 confirm that medium and large references outperform small ones, and centered references outperform off-center ones. Distance-to-edge has weaker effect once area and centeredness are controlled.
Overall, the proposed selection heuristics effectively improve one-shot performance without increasing the number of shots, providing a practical method for choosing high-value references.
A.1.4 Reference-Image Degradation
We evaluate our method under progressively degraded reference images by applying increasing levels of Gaussian blur. For efficiency, we conducted this experiment on a small representative subset, since our goal is to measure relative performance degradation rather than absolute metrics.
As shown in Figure 14, our method remains robust even under strong blur. We attribute this stability to the invariance and consistency of DINOv2 features, which remain discriminative despite significant image degradation.
Figure 14: Model performance with increasing levels of blurring in the reference image. Each column shows a reference image blurred with an increasing Gaussian kernel size, along with its corresponding DINOv2 PCA feature visualization. Blurred input images result in blurred and smoother Dinov2 features. Below each pair, we display the predicted bounding box and segmentation mask. Our method exhibits strong robustness even under heavy blur.
A.2 Few-shot Semantic Segmentation on COCO
Although our method is designed for instance segmentation, we also evaluate it on the COCO-20i Few-Shot Semantic Segmentation benchmark (Nguyen and Todorovic, 2019). The 80 COCO classes are divided into four folds (Hu et al., 2019; Wang et al., 2019; Nguyen and Todorovic, 2019), each containing 60 base and 20 novel classes. We assess performance on the 20 novel classes under strict 1-shot and 5-shot settings (Zhu et al., 2024). Results are shown in Table 5. To adapt our instance segmentation predictions to semantic segmentation, we aggregate all instances of the same class into semantic maps, enabling direct comparison with prior methods. Despite being entirely training-free, our method achieves competitive performance against fine-tuned approaches.
Methods FT 1-shot 5-shot $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean DiffewS (Zhu et al., 2024) 47.7 56.4 51.9 48.7 51.2 52.0 63.0 54.5 54.3 56.0 DiffewS-n (Zhu et al., 2024) 47.1 56.6 53.8 48.3 52.2 57.3 66.5 60.3 58.8 60.7 HSNet (Min et al., 2021) 37.2 44.1 42.4 41.3 41.2 45.9 53.0 51.8 47.1 49.5 CyCTR (Zhang et al., 2021) 38.9 43.0 39.6 39.8 40.3 41.1 48.9 45.2 47.0 45.6 VAT (Hong et al., 2022) 39.0 43.8 42.6 39.7 41.3 44.1 51.1 50.2 46.1 47.9 BAM (Lang et al., 2022) 43.4 50.6 47.5 43.4 46.2 49.3 54.2 51.6 49.6 51.2 DCAMA (Shi et al., 2022) 49.5 52.7 52.8 48.7 50.9 55.4 60.3 59.9 57.5 58.3 HDMNet (Peng et al., 2023) 43.8 55.3 51.6 49.4 50.0 50.6 61.6 55.7 56.0 56.0 Training-free 32.1 46.6 50.9 48.7 44.6 55.2 49.5 60.7 45.5 52.7
Table 5: Performance comparison of strict few-shot semantic segmentation settings (1 and 5-shot) on COCO-$20^{i}$. We aggregate instance-level predictions to allow comparison with semantic segmentation works. Previous methods results are sourced from (Zhu et al., 2024). FT refers to finetuning on novel classes.
A.3 Evaluating a VLM on COCO Few Shot Dataset
To compare with recent VLM models, we implement a joint pipeline that combines Qwen2.5-VL-7B (Wang et al., 2024) and SAM2. We run Qwen2.5-VL-7B on COCO validation set to produce category-conditioned bounding boxes, which are then directly used to prompt SAM2 to obtain segmentation masks. We evaluate both detection (nAP) and segmentation performance using the same COCO few-shot class split as our method. Table 6 shows the results obtained, highlighting the performance of our model.
Backbone models Num shots bbox segm nAP nAP50 nAP75 nAP nAP50 nAP75 DINOv2-ViT-L-14 + SAM2-L 30 36.8 54.5 38.7 34.2 54.4 36.7 Qwen2.5VL-7B-Inst + SAM2-L - 6.2 8.4 6.9 5.9 8.4 6.8
Table 6: Comparison between our DINOv2+SAM2 pipeline and the Qwen2.5-VL-7B+SAM2 baseline. Qwen-VL outputs bounding boxes that are fed to SAM2 for segmentation, while our method uses DINOv2 features with the same SAM2 decoder. Our approach achieves substantially higher detection and segmentation nAP across all metrics.
A.4 Pipeline Design
We analyze two independent design choices in our pipeline: (i) prototype construction, which defines how reference features are aggregated into category prototypes; and (ii) instance consolidation, which determines how overlapping target predictions are combined into final instances.
A.4.1 Prototype Construction
We compare (1) class-only aggregation (direct feature averaging across all class pixels) and (2) our two-step strategy, where we first compute instance-level prototypes (pixel-weighted means within each mask) and then average them to obtain the class prototype.
As shown in Table 7, the performance difference between the two strategies is negligible. This is expected: when instances have similar pixel counts, both methods produce nearly identical prototypes. The two-step version is slightly more faithful to instance structure and therefore remains our default choice.
Backbone models Aggregation strategy 10-shot 30-shot bbox segm bbox segm nAP nAP nAP nAP dinov2-sam2 class agg 35.7 33.4 36.8 34.4 dinov2-sam2 class+inst agg 35.7 33.3 36.8 34.2
Table 7: Ablation of two-step prototype aggregation strategy. We compare class-only aggregation (single-step) with two-step aggregation (instance-level pixel-weighted prototypes followed by class averaging). Results are reported for 10-shot and 30-shot settings using nAP, nAP50, and nAP75.
A.4.2 Instance Consolidation
@@
We show the incremental improvements achieved by our semantic-aware soft merging in Table 8. Other aggregation variants, such as covariance similarity, instance softmax, score decay, iterative mask refine, attention-guided global average, underperformed. Earth Mover’s Distance (EMD) was investigated in early design, but we found that its optimal-transport formulation introduced substantial computational overhead without providing clear qualitative benefits over cosine similarity.
@@
A.5 Feature Quality and Semantic Separability
A.5.1 DINOv2 Feature Separability Analysis
We compute DINOv2 features for all annotated objects in the COCO validation set (5k images) and visualize them using t-SNE. Figure 15 shows representative class pairs and triplets with varying degrees of semantic similarity.
Easily distinguishable classes (e.g., cat–dog) form well-separated clusters. In contrast, semantically similar categories (e.g., car–truck, chair–couch, wine glass–cup–vase) exhibit substantial embedding overlap. This indicates that similar-class confusion is driven by DINOv2’s feature geometry rather than prototype construction.
Improving performance in such cases requires stronger semantic disentanglement at the backbone level, which remains an open research direction.
Figure 15: t-SNE visualizations of DINOv2 object-level features for selected COCO classes. Top row: pairs of classes with clear separability (cat–dog, airplane–bear, boat–cow–orange). Bottom row: semantically similar classes with substantial embedding overlap (car–truck, chair–couch, wine glass–cup–vase). These results indicate that similar-class confusion primarily originates from DINOv2’s feature-space entanglement, limiting the effectiveness of alternative prototype-construction strategies, and setting an upper bound for our method performance.
A.5.2 Backbone comparison: DINOv2 vs. DINOv3
Feature separability. Figure 16 shows t-SNE embeddings for DINOv2-L and DINOv3-B/L/H. Classes that are not well-separated in DINOv2 remain similarly entangled in all DINOv3 variants, indicating that the semantic structure relevant for prototype-based matching does not substantially improve in DINOv3. Figure 17 provides complementary PCA visualisations of the spatial feature maps, showing that DINOv3 produces smoother and more spatially uniform features, while DINOv2 exhibits sharper local variation. These results suggest that although DINOv3 representations are denser, their object-level semantics remain comparable to DINOv2, which may limit performance improvements in our training-free setting.
Backbone size effects. Figure 18 compares model outputs across backbones. DINOv3 backbones (B/L/H) tend to assign high similarity scores more broadly across proposals.
Figure 16: t-SNE embeddings across DINO backbones. t-SNE visualisation of feature embeddings for representative classes (e.g., fork and spoon) using DINOv2-L, DINOv3-B, DINOv3-L, and DINOv3-H. Clusters that overlap in DINOv2 remain overlapping in DINOv3 variants, indicating that semantic separability does not improve across these models.
Figure 17: Feature visualisations across DINO variants. PCA projections of the spatial features for DINOv2-L and DINOv3-B/L/H.
Figure 18: Output comparison across DINOv3 backbones. Comparison of training-free outputs for DINOv2-L and DINOv3-B/L/H. In DINOv3 models, higher similarity scores are often assigned to multiple proposals, and predictions show less object-focused contrast. Ground-truth masks are shown for reference.
A.6 Implementation details
Inference Runtime Efficiency. Our training-free stages are optimised and lightweight. (1) Memory bank construction (0.1 s/img) is computed once, only requiring to encode $n$ reference images per class with DINOv2. Reference-image features are pre-cached for all following steps. (2) Semantic matching (0.0003 s/img) is a fully parallelised dot product of cosine similarities. (3) Soft-merging (0.006 s/img) uses a parallel implementation of NMS. Table 9 shows our method yields significant performance gains compared to Matcher (Liu et al., 2024c) and speeds up SAM default automatic mask generator (AMG) by $\times 3$ via efficient point sampling, faster mask filtering, and removal of unnecessary post-processing.
@@ A.6 Implementation details
Model Size and Memory Usage. We report the runtime footprint of our pipeline when using 20 classes and 10 shots per class (200 reference images). At inference time, the combined model size is 1.97 GiB, consisting of the DINOv2 encoder (1.13 GiB) and the SAM2 predictor (856 MiB). CUDA-allocated memory reaches a peak of 11.5 GiB. These measurements provide a reference for practical deployment and indicate potential avenues for integrating lighter variants (e.g., MobileSAM (Zhang et al., 2023) or LightSAM (Cheng et al., 2025)) when targeting edge devices.
A.7 Additional figures
In Figure 19, we provide extended examples comparing the PCA projections of DINOv2, DINOv3, CLIP, and PE-Spatial features, showing spatial structure and noise characteristics across backbones.
Figure 20 includes extended model output visualizations for semantic backbones, highlighting how differences in feature quality propagate to detection and segmentation predictions.
Figures 21 and 22 contain examples of similarity maps for both single-feature and aggregated-feature settings across more images and classes.
Figure 19: Extended feature comparison across semantic backbones. PCA projections of DINOv2, DINOv3, CLIP, and PE-Spatial feature maps for additional examples. These visuals highlight qualitative differences in spatial resolution, noise, and feature consistency.
Figure 20: Extended output comparison across semantic backbones. Predictions from multiple backbones on additional images, including ground-truth masks and corresponding model outputs. Differences in feature quality lead to distinct segmentation and detection behaviours. Best viewed when zoomed in.
Figure 21: Feature similarity — additional paired examples. For each reference-target image pair: the top row shows cosine similarity maps computed from a single DINOv2 patch feature (marked +++++++++++++++++), and the bottom row shows maps computed from the aggregated prototype (mask-area average). For each row we display intra-class (same image) and inter-class (target image) similarity. Paired examples demonstrate that aggregated prototypes yield more coherent, object-level similarity than single patch features.
Figure 22: Feature similarity — extended paired examples. For each reference-target image pair: the top row is single-feature similarity (patch +++++++++++++++++) and the second row is prototype similarity (mask-average). Each pair shows intra- and inter-class cosine similarity. These examples reinforce that aggregation produces more stable and spatially consistent instance-level matches.
A.8 Cross-Domain Few-Shot Object Detection
We provide additional visualisations of the six target datasets in the CD-FSOD benchmark. These datasets span diverse and challenging domains, including photorealistic, cartoon, aerial, underwater, and industrial imagery, each presenting unique distribution shifts. Despite these variations, our method achieves strong performance across all domains without any fine-tuning, demonstrating its remarkable cross-domain generalization. Figure 23 shows an overview of the results for the 6 datasets. Figures 24, 25, 27, 26, 28, 29 display more detailed results for each of the datsets. All results are shown for 5-shot setting, using 5 reference images per category.